OpenStudy (anonymous):

y= -2x^2+4x-1 y=2x-5 ???? please help

OpenStudy (anonymous):

you want the values of x and y?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

equating both we have -2x^2 +4x -1 = 2x -5 -2x^2+2x+4 =0 x^2 - x -2 =0 (x+1)(x-2)=0 x = -1, or 2 substitute in y = 2x -5 to get the values of y

OpenStudy (anonymous):

huhh ? how is that . i dont get any of thiss .

OpenStudy (anonymous):

ok so does y=-3

OpenStudy (radar):

When x=-1, y=-7 when x=2, y=-1 Were you to graph these two figures, one is a line, the other a parabola the line would cross the parabola at those points. Might be helpful to graph them and see what they look like.

OpenStudy (radar):

Oops something doesn't seem to jive. I guess I don't understand the problem.

OpenStudy (anonymous):

thanks for trying to help me thoughh !